# for Human Motion Generation

### v2

@@ Abstract
Human motion generation plays a vital role in applications such as digital humans and humanoid robot control. However, most existing approaches disregard physics constraints, leading to the frequent production of physically implausible motions with pronounced artifacts such as floating and foot sliding. In this paper, we propose Morph, a Motion-free physics optimization framework, comprising a Motion Generator and a Motion Physics Refinement module, for enhancing physical plausibility without relying on costly real-world motion data. Specifically, the Motion Generator is responsible for providing large-scale synthetic motion data, while the Motion Physics Refinement Module utilizes these synthetic data to train a motion imitator within a physics simulator, enforcing physical constraints to project the noisy motions into a physically-plausible space. These physically refined motions, in turn, are used to fine-tune the Motion Generator, further enhancing its capability. Experiments on both text-to-motion and music-to-dance generation tasks demonstrate that our framework achieves state-of-the-art motion generation quality while improving physical plausibility drastically.
@@ 1 Introduction
Accurate human motion generation is pivotal in various applications, including robotics, video gaming and virtual reality [10, 36, 45, 44, 1]. Recent advances in AI have paved the way for novel approaches to motion generation, enabling various control conditions including textual descriptions and music pieces. Mainstream approaches can be divided into three categories: conditional diffusion models [32, 42, 39, 33, 16], conditional autoregressive models [23, 13, 46, 47, 18] and generative masked modeling [7, 24]. These motion generation models have made significant progress in capturing the complex multimodal distribution of human motions, greatly improving motion generation quality.
@@
Recently, a few studies [39, 11] have attempted to enhance the physical plausibility of motion generation. For example, PhysDiff [39] and Reindiffuse [11] incorporate physical constraints into the denosing diffusion process. However, these approaches face several limitations. First, these works [39, 11] customize physics optimization operation for a specific type of generation model, i.e., text-conditional diffusion model. This customization restricts their transferability to a broader range of generation models (such as autoregressive-based models) and motion generation tasks (such as music-to-dance). Second, these works apply physics and diffusion iteratively, embedding physics optimization into multiple steps of diffusion process. The frequent execution of physics optimization significantly increases computational costs during inference. Last but most importantly, these methods require large-scale, high-quality realistic motion data to train the physics optimization mechanism. However, collecting such realistic motion data is challenging. A natural question arises: Is it possible to learn an efficient, model-agnostic physical optimizer without relying on real motion data?
@@
To achieve this, we propose a Motion-free physics optimization framework, namely Morph. As shown in Fig. 2, Morph consists of two main modules: a Motion Generator (MG) that can be any existing motion generator, and a Motion Physic Refinement (MPR) module for enhancing physical plausibility. Morph employs a two-stage optimization process. In the first stage, using large-scale synthetic noisy motion data produced by the motion generator, MPR module is optimized to project input motions into a physically-plausible space. Specifically, the MPR module designs a motion imitator that controls a character agent to mimic the given noisy motions within a physics simulator. The simulator enforces multiple physical constraints, effectively reducing artifacts such as floating and foot sliding. To further ensure the naturalness of simulated motions, MPR module introduces a motion discriminator to align the distribution of physics-refined motions with that of input motions. Feedback signals from both physics simulator and motion discriminator guide the optimization of the motion imitator via reinforcement learning. In the second stage, we observe that due to the high cost of collecting 3D motion data, existing generators typically rely on limited motion data for training, which restricts their generative capabilities. In this stage, by leveraging the trained MPR module, a large-scale set of physically plausible, high-quality motion data can be created to finetune the generator, further enhancing its ability to generate realistic motion. During inference, the finetuned motion generator and MPR module work in tandem to generate physics-plausible and high-quality motions.
@@
Through above process, Morph effectively decouples physic optimization from generation model, requiring only a single step of physical refinement and without relying on real motion data. This results in an economical, efficient, and model-agnostic physical optimizer. We evaluate our Morph framework on two motion generation tasks: text-to-motion and music-to-dance generation. Since Morph is agnostic to specific instantiation of generation models, we test it with three types of generation models, i.e., diffusion-based, autoregressive-based and generative mask modeling. Extensive experiments demonstrate that Morph achieves significant improvements in physical error metrics, while also achieving competitive generation metrics across different generators and tasks, despite not being trained on real motion data.
Figure 2: An overview of the Morph framework. Morph comprises a Motion Generator and a Motion Physics Refinement module. Morph employs a two-stage training process: Motion Physics Refinement module training and Motion Generator fine-tuning. And a Imitation Selection Operation is employed to ensure the motion quality after physics refinement.
@@ 2 Related Work
Human Motion Generation.  Motion generation is a long-history task that can be conditioned on various signals, such as text description, music and action [48, 7, 40, 8, 29, 33, 37, 2, 4, 30, 43]. Our work specifically focuses on text-to-motion [8, 40, 41, 32, 7, 24] and music-to-dance generation [33, 29, 16, 20]. Mainstream approaches can be roughly divided into three categories: diffusion-based methods [32, 3, 42, 39, 33, 16], autoregressive models [40, 17, 13, 46, 47, 18, 29] and generative masked modeling [7, 24]. For instance, MDM [32] uses a transformer-based diffusion model with conditional text representations extracted from CLIP [26]. T2M-GPT [40] designs a conditional autoregressive transformer model based on VQ-VAE [34] and GPT [25, 35]. MoMask [7] employs residual vector quantization and generative masked transformers to iteratively generate motions. For music-to-dance generation, Bailando [29] predicts discrete token sequences conditioned on music and uses an autoregressive transformer to regenerate the dance sequence. However, existing motion generation models often produce physically implausible motions, as they overlook physical laws during training. In contrast, our proposed MPR module effectively enforces physical constraints, significantly enhancing the physically plausibility of generated motions.
Physically Plausible Motion Generation. Physical plausibility refers to the degree to which generated motions adhere to physical rules, such as foot sliding, foot-ground contact, and body leaning. Generating physically-plausible motions is crucial for many real-world applications, such as animation and virtual reality. Recently, a few works have attempted to address this challenge [39, 11, 33]. For example, [39, 11] integrate physical constraints into the diffusion process, iteratively applying physics and diffusion to maintain alignment with motion data distribution while enhancing physical realism. EDGE [33] introduces auxiliary losses to align specific aspects of physical realism, specifically for generating more physically plausible dances. However, these methods tailor physic optimization to diffusion models and require real motion data, limiting their flexibility and scalability. In contrast, we propose a model-agnostic, motion-free physics optimization framework that decouples physics optimization from the generation model and learns to enhance physical plausibility without real motion data, offering a more economical and versatile solution.
@@ 3 Method
To enhance physical plausibility of motion generation without relying on real motion data, we propose a motion-free physics optimization framework, namely Morph. As illustrated in Fig. 2, Morph consists of two key modules: a Motion Generator (MG) that could be any existing pre-trained motion generation model, and a Motion Physics Refinement module (MPR) independent of Motion Generator. Morph employs a two-stage training process. In the first stage, the Motion Generator produces large-scale noisy motion data (Sec. 3.1), which is subsequently used to train the Physics Refinement module to map input motions into a physically plausible space (Sec. 3.2). In the second stage, these physics-refined motions, in turn, are used to fine-tune Motion Generator, further enhancing its capabilities (Sec 3.3). Through this two-stage training process, Morph drastically reduces physical errors while improving the overall quality of generated motions.
@@ 3.2 Physics-Based Motion Refinement
Since most existing motion generation models lack explicit physical constraints, they often produce motions with noticeable artifacts, as shown in Fig. 1. To address this, the Motion Physics Refinement (MPR) module is tasked with projecting the generated motion $\bm{\tilde{x}}^{1:L}$, which disregards the laws of physics, into a physically-plausible motion $\bm{\hat{x}}^{1:L}$. As shown in Fig. 2, the MPR module consists of three components: a motion imitator, a physics simulator and a motion discriminator. Specifically, the motion imitator controls a simulated character to mimic the input motion $\bm{\tilde{x}}^{1:L}$ within the physics simulator. The resulting motion $\bm{\hat{x}}^{1:L}$ from the physics simulator is considered physically plausible, as it adheres to the laws of physics. Additionally, the motion discriminator seeks to align the distribution of refined motions $\bm{\hat{x}}^{1:L}$ with that of input motions $\bm{\tilde{x}}^{1:L}$, further ensuring the naturalness and realism of the physics-refined motions.
@@
Motion Imitator Learning can be formulated as a Markov decision process, represented by the tuple $\left(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\right)$ of states, actions, transition dynamics and reward function. Formally, the motion imitator is described by a policy $\pi\left(\bm{a}^{l}|\bm{s}^{l}\right)$, which specifies the probability distribution of selecting an action $\bm{a}^{l}\in\mathcal{A}$ given the current state $\bm{s}^{l}\in\mathcal{S}$. The physic simulator, in turn, defines the transition dynamics $\mathcal{T}\left(\bm{s}^{l+1}|\bm{s}^{l},\bm{a}^{l}\right)$, which determines the next state $\bm{s}^{l+1}$ based on current state $\bm{s}^{l}$ and action $\bm{a}^{l}$. Specifically, starting from an initial state $\bm{s}^{1}$, a character agent acts in the physic simulator according to policy $\pi\left(\bm{a}^{l}|\bm{s}^{l}\right)$, iteratively sampling action $\bm{a}^{l}$. Then, the physic simulator, governed by the transition dynamics $\mathcal{T}\left(\bm{s}^{l+1}|\bm{s}^{l},\bm{a}^{l}\right)$, generates the next state $\bm{s}^{l+1}$, from which the simulated pose $\bm{\hat{x}}^{l+1}$ is derived. By running the policy for $L$ steps, we can obtain the simulated motion sequence $\bm{\hat{x}}^{1:L}$. In implementation, we use reinforcement learning to train the motion imitator $\pi$, where a reward is assigned based on how well the simulated motion $\bm{\hat{x}}^{1:L}$ aligns with input motion $\bm{\tilde{x}}^{1:L}$. In the following, we elaborate on the design of states, rewards, policy and actions.
States. The simulation state $\bm{s}^{l}$ consists of the input motion’s next pose, along with the differences between the input next pose and the current simulated pose across multiple aspects, including joint rotation, position, velocity and angular velocity. This difference information informs the policy of the pose residuals that require compensation, enabling the system to better align the simulated motion with the input motion. Formally, the state is defined as $\bm{s}^{l}=\left[\bm{\tilde{\theta}}^{l+1},\bm{\tilde{p}}^{l+1},\bm{\tilde{\theta}}^{l+1}-\bm{\hat{\theta}}^{l},\bm{\tilde{p}}^{l+1}-\bm{\hat{p}}^{l},\bm{\tilde{v}}^{l+1}-\bm{\hat{v}}^{l},\bm{\tilde{\omega}}^{l}-\bm{\hat{\omega}}^{l}\right]$. Here, $\bm{\theta}$, $\bm{p}$, $\bm{v}$ and $\bm{\omega}$ represent the joint rotation, position, joint velocity, and angular velocity, respectively, $\widetilde{\left[\bm{\cdot}\right]}/\widehat{\left[\bm{\cdot}\right]}$ denote the quantities of the input///simulated motions, and $\left[\bm{\cdot}\right]^{l}$ denotes the quantities at the $l^{th}$ timestep.
@@
Rewards. The reward function is designed to encourage the simulated motion to match the input motion. At each timestep $l$, the reward $r^{l}$ consists of a mimic reward $r^{l}_{\mathrm{m}}$, a energy penalty $r^{l}_{\mathrm{e}}$ and an adversarial reward $r^{l}_{\mathrm{a}}$, formulating as $r^{l}=r^{l}_{\mathrm{m}}+r^{l}_{\mathrm{e}}+r^{l}_{\mathrm{a}}$. The mimic reward $r^{l}_{\mathrm{m}}$ measures the difference between the simulated pose and input pose across multiple aspects, including joint rotation $\bm{\theta}$, position $\bm{p}$, joint velocity $\bm{v}$ and angular velocity $\bm{\omega}$, as follows:
@@
[…]
The energy penalty $r^{l}_{\mathrm{e}}$ regulates the policy by discouraging high-frequency foot jitter, a common issue in policy trained without external forces [5]. This energy penalty is computed as:
@@
The adversarial reward $r^{l}_{\mathrm{a}}$ encourages the distribution of simulated motions to align that of input motions through a motion discriminator. Given the state of current simulated pose, denoted as $\bm{\hat{s}}_{\mathrm{a}}^{l}=[\bm{\hat{\theta}}^{l},\bm{\hat{p}}^{l},\bm{\hat{v}}^{l},\bm{\hat{\omega}}^{l}]$, the adversarial reward is calculated as:
$r^{l}_{\mathrm{a}}=-\log\left(1-D_{\psi}\left(\bm{\hat{s}}_{\mathrm{a}}^{l}\right)\right).$ (4)
where $D_{\psi}$ represents the motion discriminator with parameters $\psi$.
Loss. Given the reward $r^{l}$ at each timestep $l$, we train the motion imitator using reinforcement learning (RL). The objective is to maximize the excepted reward, effectively enabling the imitator to closely mimic the input motions. To achieve this, we adopt a standard RL algorithm Proximal Policy Optimization (PPO) [28] to solve for the optimal imitator policy. The objective function is defined as:
$\mathcal{L}_{\text{PPO}}\left(\phi\right)=\mathbb{E}\left[\min\left(\eta(\phi)r^{l},\operatorname{clip}\left(\eta(\phi),1-\gamma,1+\gamma\right)r^{l}\right)\right].$ (5)
Here, $\eta(\phi)$ is the probability ratio $\frac{\pi_{\phi}\left(\bm{a}^{l}\mid\bm{s}^{l}\right)}{\pi_{\phi_{\mathrm{old}}}\left(\bm{a}^{l}\mid\bm{s}^{l}\right)}$, where $\phi$ and $\phi_{\mathrm{old}}$ refer to the parameters of the motion imitator after and before the policy update, $\operatorname{clip}$ is an operation that restricts $\eta(\phi)$ within range $\left[1-\gamma,1+\gamma\right]$ with a hyperparameter $\gamma$.
Motion Discriminator Learning.
We observe that relying solely on physics simulator often produces motions that feel overly mechanical and unnatural. To improve the naturalness of simulated motions, following [22], we introduce a motion discriminator to align the distribution of simulated motions with that of the input motions. Specifically, the motion discriminator $D_{\psi}$ learns to distinguish between the state sampled from input pose ($\bm{\tilde{s}}_{\mathrm{a}}^{l}$) and that of simulated pose ($\bm{\hat{s}}_{\mathrm{a}}^{l}$), with the objective function as:
$\mathcal{L}_{\text{Dis}}\left(\psi\right)$ $=-\mathbb{E}\left[\log\left(D_{\psi}\left(\bm{\tilde{s}}_{\mathrm{a}}^{l}\right)\right)\right]-\mathbb{E}\left[\log\left(1-D_{\psi}\left({\bm{\hat{s}}_{\mathrm{a}}^{l}}\right)\right)\right],$ (6)
where $\psi$ denotes the parameters of the motion distribution.
@@
During training, we employ reference state initialization [21], where a starting point is randomly selected from a motion clip for imitation. We observe that the character frequently falls in the early stages of training. To improve training efficiency and accelerate convergence, we adopt an early stopping strategy. Specifically, we terminate the episode when the MPJPE (Mean Per Joint Position Error) between the state of input pose and that of simulated pose exceeds $0.5$ meter.
@@
As training progresses, the motion imitator gradually learns to imitate simple motion sequences. However, more challenging examples in the large-scale motion dataset may be overlooked, limiting the model’s ability to handle difficult samples. To address this, we implement a Hard Negative Mining process that identifies motions where the physical simulator fails to imitate as hard samples. Specifically, a dynamic weight is assigned to each motion sequence in the input data, doubling whenever imitation fails. This process progressively increases the focus on challenging samples, guiding the imitator to effectively learn from difficult examples.

### v3

@@ Abstract
Human motion generation has been widely studied due to its crucial role in areas such as digital humans and humanoid robot control. However, many current motion generation approaches disregard physics constraints, frequently resulting in physically implausible motions with pronounced artifacts such as floating and foot sliding. Meanwhile, training an effective motion physics optimizer with noisy motion data remains largely unexplored. In this paper, we propose Morph, a Motion-Free physics optimization framework, consisting of a Motion Generator and a Motion Physics Refinement module, for enhancing physical plausibility without relying on expensive real-world motion data. Specifically, the motion generator is responsible for providing large-scale synthetic, noisy motion data, while the motion physics refinement module utilizes these synthetic data to learn a motion imitator within a physics simulator, enforcing physical constraints to project the noisy motions into a physically-plausible space. Additionally, we introduce a prior reward module to enhance the stability of the physics optimization process and generate smoother and more stable motions. These physically refined motions are then used to fine-tune the motion generator, further enhancing its capability. This collaborative training paradigm enables mutual enhancement between the motion generator and the motion physics refinement module, significantly improving practicality and robustness in real-world applications. Experiments on both text-to-motion and music-to-dance generation tasks demonstrate that our framework achieves state-of-the-art motion quality while improving physical plausibility drastically.
Figure 1: Examples of physical inconsistencies in generations.
@@ 1 Introduction
Accurate human motion synthesis plays a crucial role in applications such as robotics, video gaming, and virtual reality [9, 38, 47, 46, 1]. Recent advances in AI have enabled motion generation under various control conditions, including textual descriptions and music inputs. Current approaches can be categorized into three main paradigms: conditional diffusion models [34, 44, 41, 35, 15], conditional autoregressive models [26, 12, 48, 49, 19], and generative masked modeling [6, 27]. These models have significantly improved the quality of generated motion by better capturing the complex multimodal distribution of human movement.
@@
Recently, a few studies [41, 10] have attempted to enhance the physical plausibility of motion generation, making the generated motions more realistic. For example, PhysDiff [41] and Reindiffuse [10] incorporate physical constraints into the denoising diffusion process. However, these approaches face two main limitations. On the one hand, although these works [41, 10] introduce promising approaches for motion physics optimization, their generalizability and transferability across different types of motion generation models and tasks have not been fully demonstrated or explored. On the other hand, these works apply physics and diffusion iteratively, embedding physics optimization into multiple steps of diffusion process. The frequent execution of physics optimization significantly increases computational costs during inference. Additionally, prior studies [22, 21, 20] have explored physics-based human motion imitation using a physics simulator, typically relying on large-scale real 3D motion data as priors to enhance accuracy and naturalness in imitation learning. However, collecting large-scale realistic motion data is challenging. A natural question arises: Is it possible to learn an efficient, model-agnostic physical optimizer without relying on real motion data?
@@
To achieve this, we propose a Motion-free physics optimization framework, namely Morph. As shown in Fig. 2, Morph consists of two main modules: a Motion Generator (MG) that can be any existing motion generator, and a Motion Physic Refinement (MPR) module for enhancing physical plausibility. Morph employs a two-stage optimization process. In the first stage, using large-scale noisy motion data produced by the motion generator, MPR module is optimized to project input motions into a physically-plausible space. Specifically, MPR module designs a motion imitator that controls a character agent to mimic the given noisy motions within a physics simulator. The simulator enforces multiple physical constraints, effectively reducing artifacts such as floating and foot sliding. To further ensure the naturalness of simulated motions, MPR module introduces a prior reward module to align the distribution of physics-refined motions with that of input motions. This prior reward module can provide smoother and more stable rewards. Feedback signals from both physics simulator and prior reward module guide the optimization of the motion imitator via reinforcement learning. In the second stage, we address the limitations of pretrained motion generators, which may not be trained on data that adheres to physical constraints and thus produce suboptimal results. To this end, we finetune the motion generator using the physic-refined motion data produced by MPR module, further enhancing its generation capability. During inference, the fine-tuned motion generator and MPR module work in tandem to generate physics-plausible and high-quality motions.
@@
By fostering the synergy between motion generation and motion physics refinement, morph provides a scalable, efficient, and generalizable solution to realistic human motion synthesis. We evaluate our morph framework on text-to-motion and music-to-dance tasks, using diffusion-based, autoregressive, and generative mask models. Extensive experiments demonstrate that morph achieves significant reduces physical errors, while maintaining competitive motion quality across different generators and tasks, despite not being trained on real motion data.
Our contributions can be summarized as follows:
  • We introduce Morph, a generalizable framework that enhances physical plausibility and generation quality across various motion generation models and tasks. Also, a prior reward module is designed to accelerate imitation learning and improve motion realism in physics simulation.
  • We design a collaborative training paradigm, where the Motion Generator provides large-scale synthetic motion data to enhance the Motion Physics Refinement module, which, in turn, refines motion quality by feeding back physics-refined data.
  • Extensive qualitative and quantitative experiments demonstrate that Morph achieves competitive physical plausibility and generation performance.
@@ 2 Related Work
Human Motion Generation. Motion generation is a long-history task that can be conditioned on various signals, such as text description, music and action [50, 6, 42, 7, 31, 35, 39, 16, 17, 2, 4, 32, 45]. Our work specifically focuses on text-to-motion [7, 42, 43, 34, 6, 27] and music-to-dance generation [35, 31, 15, 23]. Mainstream approaches can be roughly divided into three categories: diffusion-based methods [34, 3, 44, 41, 35, 15], autoregressive models [42, 18, 12, 48, 49, 19, 31] and generative masked modeling [6, 27]. For instance, MDM [34] uses a transformer-based diffusion model with conditional text representations extracted from CLIP [29]. T2M-GPT [42] designs a conditional autoregressive transformer model based on VQ-VAE [36] and GPT [28, 37]. MoMask [6] employs residual vector quantization and generative masked transformers to iteratively generate motions. For music-to-dance generation, Bailando [31] predicts discrete token sequences conditioned on music and uses an autoregressive transformer to regenerate the dance sequence. However, existing motion generation models often produce physically implausible motions, as they overlook physical laws during training. In contrast, our proposed MPR module effectively enforces physical constraints, significantly enhancing the physically plausibility of generated motions.
Figure 2: An overview of the Morph framework. Morph comprises a Motion Generator and a Motion Physics Refinement module. Morph employs a two-stage training process: Motion Physics Refinement module training and Motion Generator fine-tuning. And a Imitation Selection Operation is employed to ensure the motion quality after physics refinement. The solid curved arrows on the left and right (in orange and green) represent the iterative, collaborative optimization between Stage 1 and Stage 2.
Physically Plausible Motion Generation. Physical plausibility refers to the degree to which generated motions adhere to physical rules, such as foot sliding, foot-ground contact, and body leaning. Generating physically-plausible motions is crucial for many real-world applications, such as animation and virtual reality. Recently, a few works have attempted to address this challenge [41, 10, 35]. For example, [41, 10] integrate physical constraints into the diffusion process, iteratively applying physics and diffusion to maintain alignment with motion data distribution while enhancing physical realism. EDGE [35] introduces auxiliary losses to align specific aspects of physical realism, specifically for generating more physically plausible dances. However, these methods rely on real motion data and have not fully demonstrated generalizability across different motion generation models and tasks, limiting their flexibility and scalability.
@@ 3 Method
To enhance physical plausibility of motion generation without relying on real motion data, we propose a motion-free physics optimization framework, namely morph. As illustrated in Fig. 2, morph consists of two key modules: a Motion Generator (MG) that could be any existing pre-trained motion generation model, and a Motion Physics Refinement module (MPR) independent of Motion Generator. Morph employs a two-stage training process. In the first stage, the Motion Generator produces large-scale noisy motion data (Sec. 3.1), which is subsequently used to train the Physics Refinement module to map input motions into a physically plausible space (Sec. 3.2). In the second stage, these physics-refined motions, in turn, are used to fine-tune Motion Generator, further enhancing its capabilities (Sec 3.3). Through these two-stage training process, morph drastically reduces physical errors while improving the overall quality of generated motions.
@@ 3.2 Physics-Based Motion Refinement
Since most existing motion generation models lack explicit physical constraints, they often produce motions with noticeable artifacts, as shown in Fig. 1. To address this, the Motion Physics Refinement (MPR) module is tasked with projecting the generated motion $\bm{\tilde{x}}^{1:L}$, which disregards the laws of physics, into a physically-plausible motion $\bm{\hat{x}}^{1:L}$. As shown in Fig. 2, the MPR module consists of three components: a motion imitator, a physics simulator and a motion prior reward model. Specifically, the motion imitator controls a simulated character to mimic the input motion $\bm{\tilde{x}}^{1:L}$ within the physics simulator. The resulting motion $\bm{\hat{x}}^{1:L}$ from the physics simulator is considered physically plausible, as it adheres to the laws of physics. Additionally, the prior reward model penalize the non-human motions further ensuring the naturalness and realism of the physics-refined motions.
@@
Motion Imitator Learning can be formulated as a Markov decision process, represented by the tuple $\left(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\right)$ of states, actions, transition dynamics and reward function. Formally, the motion imitator is described by a policy $\pi\left(\bm{a}^{l}|\bm{s}^{l}\right)$, which specifies the probability distribution of selecting an action $\bm{a}^{l}\in\mathcal{A}$ given the current state $\bm{s}^{l}\in\mathcal{S}$. The physic simulator, in turn, defines the transition dynamics $\mathcal{T}\left(\bm{s}^{l+1}|\bm{s}^{l},\bm{a}^{l}\right)$, which determines the next state $\bm{s}^{l+1}$ based on current state $\bm{s}^{l}$ and action $\bm{a}^{l}$. Specifically, starting from an initial state $\bm{s}^{1}$, a character agent acts in the physic simulator according to policy $\pi\left(\bm{a}^{l}|\bm{s}^{l}\right)$, iteratively sampling action $\bm{a}^{l}$. Then, the physic simulator, governed by the transition dynamics $\mathcal{T}\left(\bm{s}^{l+1}|\bm{s}^{l},\bm{a}^{l}\right)$, generates the next state $\bm{s}^{l+1}$, from which the simulated pose $\bm{\hat{x}}^{l+1}$ is derived. By running the policy for $L$ steps, we can obtain the simulated motion sequence $\bm{\hat{x}}^{1:L}$. In implementation, we adopt a standard RL algorithm Proximal Policy Optimization (PPO) to train the motion imitator $\pi$, where a reward is assigned based on how well the simulated motion $\bm{\hat{x}}^{1:L}$ aligns with input motion $\bm{\tilde{x}}^{1:L}$. In the following, we elaborate on the design of states, rewards, policy and actions.
States. The simulation state $\bm{s}^{l}$ consists of the input motion’s next pose, along with the differences between the input next pose and the current simulated pose across multiple aspects, including joint rotation, position, velocity and angular velocity. This difference information informs the policy of the pose residuals that require compensation, enabling the system to better align the simulated motion with the input motion. Formally, the state is defined as $\bm{s}^{l}=\left[\bm{\tilde{\theta}}^{l+1},\bm{\tilde{p}}^{l+1},\bm{\tilde{\theta}}^{l+1}-\bm{\hat{\theta}}^{l},\bm{\tilde{p}}^{l+1}-\bm{\hat{p}}^{l},\bm{\tilde{v}}^{l+1}-\bm{\hat{v}}^{l},\bm{\tilde{\omega}}^{l}-\bm{\hat{\omega}}^{l}\right]$. Here, $\bm{\theta}$, $\bm{p}$, $\bm{v}$ and $\bm{\omega}$ represent the joint rotation, position, joint velocity, and angular velocity, respectively, $^\widetilde{\left[\bm{\cdot}\right]}$$\widehat{\left[\bm{\cdot}\right]}$ denote the quantities of the input//simulated motions, and $\left[\bm{\cdot}\right]^{l}$ denotes the quantities at the $l^{th}$ timestep.
@@
Rewards. The reward function is designed to encourage the simulated motion to match the input motion. At each timestep $l$, the reward $r^{l}$ consists of a mimic reward $r^{l}_{\mathrm{m}}$ and a energy penalty $r^{l}_{\mathrm{e}}$ and a prior reward $r^{l}_{\mathrm{p}}$, formulating as $r^{l}=r^{l}_{\mathrm{m}}+r^{l}_{\mathrm{e}}+r^{l}_{\mathrm{p}}$. The mimic reward $r^{l}_{\mathrm{m}}$ includes joint rotation $\bm{\theta}$, position $\bm{p}$, joint velocity $\bm{v}$ and angular velocity $\bm{\omega}$, as follows:
@@
where $w_{\theta}$, $w_{p}$, $w_{v}$, $w_{\omega}$, are weighting factors, $\alpha_{\theta}$, $\alpha_{p}$, $\alpha_{v}$, $\alpha_{\omega}$ are scaling factors, $^\widetilde{\left[\bm{\cdot}\right]}$$\widehat{\left[\bm{\cdot}\right]}$ denote the quantities of the input//simulated pose, and $\left|\cdot\right|$ is the L1 norm.
The energy penalty is computed as:
@@
Prior Reward Module.
Relying solely on physical simulators and imitation learning often results in non-human-like actions, characterized by a mechanical feel and unrealistic rotations. Previous works [25, 21] have attempted to address these issues by employing adversarial-based methods trained on high-quality motion data to incorporate human motion knowledge, improving the naturalness of the motion. However, this adversarial-based method often suffers from training instability and produces rewards that are not smooth enough, which also prolongs the overall imitation learning training process. To address this issue, we propose a Motion VAE Reward model as a prior to constrain the motion generated by the physics simulator, making it more natural. As shown in Fig. 2, to accelerate the training speed of motion imitation learning, we first train a lightweight Motion VAE model using motion data generated by the motion generator. This allows the model to effectively learn the distribution of the generated data. The objective function for optimizing the Motion VAE model is as follows:
$\mathcal{L}(\theta,\phi;\mathbf{s})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{\Delta s})}\left[\log p_{\theta}(\mathbf{\Delta s}|\mathbf{z})\right]-D_{\text{KL}}\left(q_{\phi}(\mathbf{z}|\mathbf{\Delta s})\parallel p(\mathbf{z})\right)$ (4)
where $\theta$ and $\phi$ denote the parameters of VAE, $\Delta s$ refer to the difference in state between two consecutive time steps for reference motion. $D_{KL}(\cdot,\cdot)$ refers to the KL divergence.
As shown in Fig. 2, after the Motion VAE model is trained, the pre-trained Motion VAE model is applied to compute rewards as follows:
[…]
where $\Delta s$ and $\Delta\overline{s}$ refer to the difference in state between two consecutive time steps for simulated motion and the reconstructed simulated motion, respectively. $\left\|\cdot\right\|$ is the L1 norm.
@@
Similar to [24], a starting point is randomly selected from a motion clip for imitation. To improve training efficiency and accelerate convergence, we terminate the episode when the MPJPE (Mean Per Joint Position Error) between the state of input pose and that of simulated pose exceeds $0.5$ meter.
@@
As training progresses, the motion imitator gradually learns to imitate simple motion sequences. However, more challenging examples in the large-scale motion dataset may be overlooked, limiting the model’s ability to handle difficult samples. To address this, we implement a Hard Negative Mining process that identifies motions where the physical simulator fails to imitate as hard samples. Specifically, a dynamic weight is assigned to each motion sequence for sampling, doubling whenever imitation fails. This process progressively increases the focus on challenging samples, guiding the imitator to effectively learn from difficult examples.